Hubbard Industries just paid a common dividend, D0, of $1.90. It expects to grow at a constant rate of 3% per year. If investors require a 11% return on equity, what is the current price of Hubbard's common stock? Do not round intermediate calculations. Round your answer to the nearest cent.
Answer #1

D0 = 1.90, g = 3%, ke = 11%

D1 = D0*(1+g) = 1.90*(1+0.03) = 1.957

formula : P0 = D1/(ke-g)

P0 = D1/(ke-g) = 1.957/(0.11 - 0.03) = 24.46

ANSWER : 24.46 (Thumbs up please)